# µOpTime: Statically Reducing the Execution Time of Microbenchmark Suites Using Stability Metrics

### v1

@@ 5. Discussion
In our experiments and also in the data sets used, each microbenchmark was repeated many times to ensure reliable results. For single specific microbenchmarks, however, this full execution configuration might not repeat the microbenchmark “often enough” to yield reliable results. This might be because it would require even more suite repetitions or iterations, making the suite execution time even longer and, therefore, infeasible to execute. We see this especially when applying µOpTime: Ideally, running the full execution configuration should repeat each microbenchmark more often than necessary to ensure relevant results. With our optimization, there should then be a reduction in each microbenchmark, which is not always the case. In the ten Java projects of data set 2 we examined, the execution configurations of 1,414 out of 3,969 microbenchmarks (35.63%) were not reduced, when RCIW3 was used as a stability metric, and warmup is considered. Thus, stability measure and configuration must be aligned, and in doubt, more repetitions must be added to the full execution configuration. Additionally, as the stability of a microbenchmark might change over time, a full execution configuration might become insufficient over time. This problem can only be caught, if the full execution is repeated. We evaluated, how well µOpTime performs in CI/CD pipeline in Section 4.5 compared to the full execution, but we do not consider the problem of the full execution becoming insufficient itself. When collecting data set 1, we increased the parameter seconds per iteration from 1 second in previous work to 5 seconds for our data collection to mitigate this problem, as the experiments ran in a cloud environment.
@@ 7. Related Work
A third approach is to apply test case prioritization techniques to the benchmark area (Rothermel et al., 1999; Mostafa et al., 2017; Laaber et al., 2021b, 2022). Running critical and relevant microbenchmarks first that already detect any degradation in the first iterations makes the execution of further benchmarks unnecessary and delays the result. As µOpTime uses RMIT execution of microbenchmark suites and always explores the whole parameter space of measurements, a prioritization of individual microbenchmarks only makes sense when using the reduced execution configurations in regular CI/CD pipelines; however, finding ways to leverage prioritization and µOpTime together is desirable and subject to future research.

### v2

@@ 5. Discussion
In our experiments and also in the data sets used, each microbenchmark was repeated many times to ensure reliable results. For single specific microbenchmarks, however, this full execution configuration might not repeat the microbenchmark “often enough” to yield reliable results. This might be because it would require even more suite repetitions or iterations, making the suite execution time even longer and, therefore, infeasible to execute. We see this especially when applying µOpTime: Ideally, running the full execution configuration should repeat each microbenchmark more often than necessary to ensure relevant results. With our optimization, there should then be a reduction in each microbenchmark, which is not always the case. In the ten Java projects of data set 2 we examined, the execution configurations of 1,414 out of 3,969 microbenchmarks (35.63%) were not reduced, when RCIW3 was used as a stability metric, and warmup is considered. Thus, stability measure and configuration must be aligned, and in doubt, more repetitions must be added to the full execution configuration. Additionally, as the stability of a microbenchmark might change over time, a full execution configuration might become insufficient over time. This problem can only be caught, if the full execution is repeated. We evaluated, how well µOpTime performs in CI/CD pipeline in Section˜4.5 compared to the full execution, but we do not consider the problem of the full execution becoming insufficient itself. When collecting data set 1, we increased the parameter seconds per iteration from 1 second in previous work to 5 seconds for our data collection to mitigate this problem, as the experiments ran in a cloud environment.
@@ 7. Related Work
A third approach is to apply test case prioritization techniques to the benchmark area (Rothermel et al., 1999; Mostafa et al., 2017; Laaber et al., 2021b, 2024). Running critical and relevant microbenchmarks first that already detect any degradation in the first iterations makes the execution of further benchmarks unnecessary and delays the result. As µOpTime uses RMIT execution of microbenchmark suites and always explores the whole parameter space of measurements, a prioritization of individual microbenchmarks only makes sense when using the reduced execution configurations in regular CI/CD pipelines; however, finding ways to leverage prioritization and µOpTime together is desirable and subject to future research.